OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4).

OpenStudy (anonymous):

Answer Choices: y = 4/3x + 8 y = 4/3x + 3 y = −3/4x + 8 y = −3/4x + 3

OpenStudy (anonymous):

change the original equation so that you get the gradient 4x - 3y = -12 -3y = -12 -4x y = 4 + 4/3x so the gradient is 4/3 the second equation will become y = 4/3x + c now to find c, substitute the coordinates into the equation 4 = -12/3 + c 4 = -4 + c c = 8 so the equation should be y = 4/3x +8

OpenStudy (anonymous):

Can you help me with one more.?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2).

OpenStudy (anonymous):

this question is very similiar to the previous one change the equation such that we can get the gradient -3y = -2x +3 y = 2/3x -1 so the gradient is 2/3 well in this case for perpendicular lines, the product of the gradient is -1. divide -1 by 2/3 and we get -3/2 the equation we get will be y = -3/2x + c to find c, substitute the coordinates into the equation 2 = 24/2 + c 2 = 12 + c c = -10 so the equation is y = -3/2x -10

OpenStudy (anonymous):

Thank you so much (: